OpenStudy (anonymous):

Solve method by the eliminated method.... 5x+5y=-7 7x-3y=11

OpenStudy (anonymous):

order pair or simplified fraction

OpenStudy (2bornot2b):

Would you like your solution here?

OpenStudy (2bornot2b):

\[x = 17/25, y = -52/25 \]

OpenStudy (anonymous):

so -52, 25 is the order pair

OpenStudy (2bornot2b):

No , your ordered pairs are \[(17/25,−52/25)\]

OpenStudy (anonymous):

that said incorrect

OpenStudy (2bornot2b):

who said thats incorrect?

OpenStudy (anonymous):

my mathlab access say correct if it is and incorrect if it is... there are also the answerr infiniately many answer and no solution

OpenStudy (anonymous):

Sorry I forgot to put ) in there and it said wrong

OpenStudy (2bornot2b):

I am not gettin you, is it correct or not?

OpenStudy (anonymous):

yes sorry it was correct my type o made it wrong